A box is pulled across a level floor a distance of 0.1 km. Given that 2.0 kJ of work was done in overcoming friction, what was the average friction force?


1
Expert's answer
2022-05-18T15:28:09-0400
"W=Fd\cos\theta"

"F=\frac{W}{d}=\frac{2000\: \rm J}{100\: \rm m}=20\:\rm N"